OpenStudy (anonymous):

Is the system consistent or inconsistent? y=-1/2x-1 and y=3x+3

OpenStudy (anonymous):

which is the first equation? its not completely clear \[y =\frac{-1}{2x-1} \text{ or } y = \frac{-1}{2x} -1\]

OpenStudy (anonymous):

Honestly, I don't know. It's typed out exactly how I posted it.

OpenStudy (anonymous):

My best guess would be the second equation that you posted.

OpenStudy (anonymous):

neither of them are inconsistent

OpenStudy (anonymous):

i'll explain how in just a minute

OpenStudy (anonymous):

ok, so inconsistent means that there are no values of y or x that satisfy those equations, in effect we cant solve them because we will end up with something silly like 0 = 3

OpenStudy (anonymous):

1/2 x (half x) is also possible...

OpenStudy (anonymous):

where is this question from? its not very clear what it means

OpenStudy (anonymous):

From my homework.

OpenStudy (anonymous):

what kind of level maths is it if you dont mind me asking? i could guess at what the equations are because they are of different difficulty

OpenStudy (anonymous):

Algebra 1

OpenStudy (anonymous):

hmm im not sure... :/

OpenStudy (anonymous):

usually we talk about linear equations being inconsistent

OpenStudy (anonymous):

The answer was consistent, haha. I got it right.

OpenStudy (anonymous):

so maybe we mean: \[y=-\frac{1}{2}x-1\]

OpenStudy (radar):

-1/2 x - 1 = 3x + 3 -3 1/2 x = 4 x= - 4/3 1/2 -4 * 2/7=-8/7 x= - 8/7 y= 3x + 3 y = 3 (-8/7) +3 y =-24/7 + 3 y= -3/7 Do you have an option similar to those values?

OpenStudy (radar):

That system is consistent (if that is what it is)